If the square of 2 less than an integer is 16, find the number(s). I simply do not understand the questio, much less know how to answer it. Help would be greatly appreciated. Thanks soooo much!!
Answer by edjones(8007)

let x= an integer.
(x-2)^2=16
x-2=+-4 square root of both sides.
x=2+4=6, x=2-4=-2 add 2 to both sides.
Ed